# Establishing Provenance Before Coding: Traditional and Next-Gen Signing

### v1

@@ CONCLUSIONS
Finally, we remind the reader that provenance — even cryptographically-assured provenance using signatures — is not a guarantee of correctness nor security. Provenance information reduces the risk of certain classes of attacks, such as man-in-the-middle substitutions. Provenance information also improves the engineer’s ability to estimate the quality of the components they rely on, e.g., based on the reputation of the supplier. But guaranteeing the correctness and security of a component (not to mention the resulting system) is a task for formal methods, which is another endeavor for computing.
@@ ACKNOWLEDGMENTS
The authors thank Zach Steindler of GitHub for his feedback on our estimates of npm provenance adoption. This work was supported by the US National Science Foundation under award #2229703, and by funds from Google and Cisco.

### v2

@@ CONCLUSIONS
Finally, we caution the reader that provenance — even cryptographically-assured provenance using signatures — does not guarantee correctness nor security. Provenance information reduces the risk of certain classes of attacks, such as man-in-the-middle substitutions. Provenance information also improves the engineer’s ability to estimate the quality of the components they rely on, e.g., based on the reputation of the supplier. But guaranteeing the correctness and security of a component (not to mention the resulting system) is a task for formal methods, which is another endeavor for computing.
@@ ACKNOWLEDGMENTS
The authors thank Zach Steindler of GitHub for his feedback on our estimates of npm provenance adoption, and Chinenye Okafor for her input on the design of next-gen signing systems. This work was supported by the US National Science Foundation under award #2229703, and by funds from Google and Cisco.